# Efficient and Flexible Method for Reducing Moderate-size Deep Neural Networks with Condensation

### v1

@@ 3.2.3 Reduction Result
In the CIFAR10 image classification task, we can quickly measure the performance of neural networks using the accuracy of the validation set, thus, we adopt automatic condensation reduction. The total condensation reduction training process is shown in the figure 6(a) and figure 6(b). From the start of training to the beginning of the first major reduction, it took 293 steps, at which point the model reached a validation set accuracy of 88.16%. The major reduction was completed in 338 steps, reducing from the 2nd layer to the 17th layer, totaling 16 reductions. Initially, the main reduction criterion was set at 0.88 and the layer-by-layer reduction criterion at 0.84; by the end, these were adjusted to 0.87837 and 0.83837, respectively. During this major reduction, the condensation threshold for each layer was maintained at the original value of 0.88080. The original model’s parameter count was reduced from 2,236,682 to 1,143,421, meaning the reduced model’s parameters accounted for only 51.1% of the original model. As shown in the figure 7, throughout the major reduction process, most layers underwent significant reductions, with reduction ratios around 50%. The greatest reduction was in the 17th layer, with a reduction ratio of 19.69% and a single-layer parameter reduction of 380,103. Generally, the model exhibited minor performance deviations after each layer’s reduction, with post-reduction, untrained validation set accuracies mostly above 70%. After significant reductions, accuracy could reach over 80% with just one step of training per layer. The observed reduction in parameter count validates the prevalence of condensation phenomena in convolutional neural networks, with a significant number of neurons condensing around a condensation threshold of approximately 0.88. The minor shifts resulting from reductions confirm that the condensation reduction method extended from fully connected to convolutional neural networks is reasonable and effective. The ability to rapidly restore model performance post-reduction suggests that the critical points achieved by the original model likely coincide with those of the reduced subnetworks. At step 385, 47 steps post-reduction, the reduced model’s validation set accuracy exceeded 87%. Considering model deployment objectives, it’s only necessary to proceed with an additional 45 steps of reduction training after meeting deployment requirements to reduce nearly half of the parameter count. An additional 47 steps of training can yield a reduced model with negligible performance differences from the original, offering substantial practical value.
@@
Figure 6: Training loss and accuracy with parameter nums of CNN model. Subfigures (a) and (b) respectively show the relationship between model parameter nums and accuracy, and training error. It is clearly observable that as the number of parameters decreases, accuracy tends to decrease and training error tends to increase. However, both metrics quickly return to their original levels.
Figure 7: First time of main redcution. The red reduction ratio indicates the proportion of the layer’s parameters after reduction relative to the original number of parameters; a smaller ratio suggests a greater reduction. The blue section represents the accuracy of the model after reduction. ”Acc 0 epoch” denotes the accuracy of the model immediately after reduction but before any training, whereas ”Acc 1 epoch” represents the accuracy after the model has been reduced and trained for one step.

### v2

@@ 3.2.3 Reduction Result
In the CIFAR10 image classification task, we can quickly measure the performance of neural networks using the accuracy of the validation set, thus, we adopt automatic condensation reduction. Please refer to Table 3 for a comparison of the reduction processes. The total condensation reduction training process is shown in the figure 6(a) and figure 6(b). From the start of training to the beginning of the first major reduction, it took 293 steps, at which point the model reached a validation set accuracy of 88.16%. The major reduction was completed in 338 steps, reducing from the 2nd layer to the 17th layer, totaling 16 reductions. Initially, the main reduction criterion was set at 0.88 and the layer-by-layer reduction criterion at 0.84; by the end, these were adjusted to 0.87837 and 0.83837, respectively. During this major reduction, the condensation threshold for each layer was maintained at the original value of 0.88080. The original model’s parameter count was reduced from 2,236,682 to 1,143,421, meaning the reduced model’s parameters accounted for only 51.1% of the original model. As shown in the figure 7, throughout the major reduction process, most layers underwent significant reductions, with reduction ratios around 50%. The greatest reduction was in the 17th layer, with a reduction ratio of 19.69% and a single-layer parameter reduction of 380,103. Generally, the model exhibited minor performance deviations after each layer’s reduction, with post-reduction, untrained validation set accuracies mostly above 70%. After significant reductions, accuracy could reach over 80% with just one step of training per layer. The observed reduction in parameter count validates the prevalence of condensation phenomena in convolutional neural networks, with a significant number of neurons condensing around a condensation threshold of approximately 0.88. The minor shifts resulting from reductions confirm that the condensation reduction method extended from fully connected to convolutional neural networks is reasonable and effective. The ability to rapidly restore model performance post-reduction suggests that the critical points achieved by the original model likely coincide with those of the reduced subnetworks. At step 385, 47 steps post-reduction, the reduced model’s validation set accuracy exceeded 87%. Considering model deployment objectives, it’s only necessary to proceed with an additional 45 steps of reduction training after meeting deployment requirements to reduce nearly half of the parameter count. An additional 47 steps of training can yield a reduced model with negligible performance differences from the original, offering substantial practical value.
Table 3: Comparison table of reduction process of CNN. ”Parameter” represents the number of parameters in the neural network. Reduction rate represents the ratio of the current network parameter count to the original network parameter count. A lower reduction rate indicates a smaller network relative to the original, demonstrating a more effective reduction. Accuracy refers to the accuracy on the test set. From the table, it is clear that even with significant changes in the reduction rate, the accuracy only fluctuates slightly, showcasing the coalescence reduction algorithm’s effectiveness in reducing the network size while maintaining model performance.
Model Parameter Reduction ratio Accuracy
Original 2,236,682 100% 88.16%
1st reduced 1,143,421 51.12% 88.01%
2nd reduced 910,370 40.70% 88.00%
3rd reduced 827,032 36.98% 87.24%
4th reduced 776,011 34.69% 86.36%
5th reduced 735,527 32.88% 85.88%
6th reduced 704,004 31.48% 85.80%
7th reduced 669,228 29.92% 83.88%
Final reduced 258,212 11.54% 83.21%
@@
Figure 6: Training loss and accuracy with parameter nums of CNN model. CPN stands for Current Parameter Number, OPN stands for Original Parameter Number, and CPN/OPN represents their ratio. Each red dot in the figure represents a single-layer reduction. Subfigures (a) and (b), respectively, show the relationship between model parameter nums and accuracy and training error. It is clearly observable that as the number of parameters decreases, accuracy tends to decrease and training error tends to increase. However, both metrics quickly return to their original levels.
Figure 7: First time of main redcution. The red reduction ratio indicates the proportion of the layer’s parameters after reduction relative to the original number of parameters; a smaller ratio suggests a greater reduction. The green “Acc 0 epoch” represents the accuracy of the network after reduction without any further training, while the blue “Acc 1 epoch” represents the accuracy of the network after reduction with only one step of training.